a tank holds 80 liters of a chemical solution. currently, the solution has a strenght of 30%. how much of this solution must be drained and replaced with a 70% solution to get a strenght of 40%?
Answer by ankor@dixie-net.com(22740)

a tank holds 80 liters of a chemical solution. currently, the solution has a strenght of 30%. how much of this solution must be drained and replaced with a 70% solution to get a strenght of 40%?
:
Let x = amt to be drained and it's also the amt to be added
:
The final solution will be still be 80 liters
:
30% Solution after x is drained: (80-x)
:
70% Solution added = x
:
A simple equation:
.30(80-x) + .70x = .40(80)
24 - .3x + .7x = 32
.4x = 32 - 24
.4x = 8
x = 8/.4
x = 20 gallons drained and replaced
:
:
Check:
.30(60) + .7(20) = .4(80)
18 + 14 = 32